how much money do i need to net EXACTLY $25,000 in 8 years at 5%interest compounded annually? Carrying out three decimal points then round to nearest $
Answer by Alan3354(69443)

how much money do i need to net EXACTLY $25,000 in 8 years at 5%interest compounded annually?
--------------------
Interest = AMT*(1.05)^n - AMT = 25000
AMT*(-1 + 1.05^8) = 25000
AMT = 25000/(-1 +1.05^8)
AMT =~ $52,360.9068